Two angles A and B are complementary. If angle A is 20 degrees larger than angle B what is the measure of angle A?
Answer by stanbon(75887)

Two angles A and B are complementary. If angle A is 20 degrees larger than angle B what is the measure of angle A?
------------
Equations:
A + B = 90
A = B+20
------------------------
Substitute for "A" and solve for "B":
B+20+B = 90
2B = 70
B = 35
----
then A = 35+20 = 55 degrees
=================================
Cheers,
Stan H.